Verify that the equations
i) z =sqrt (2x + a )+ sqrt(2y + b) and
ii)z^2+u=2(1+l ^x)(x+ly)
are both complete integrals of the PDEz=1/p+1/q .
Also show that the complete integral
(ii) is the envelope of one parameter sub-system obtained by taking b=-a/l -μ/1+l in the
solution (i)
1
Expert's answer
2020-03-30T12:35:24-0400

Question (1i)



"z=\sqrt{2x+a}+\sqrt{2y+b}\longrightarrow\\[0.3cm]
p=\frac{\partial z}{\partial x}=\frac{2}{2\sqrt{2x+a}}\equiv\frac{1}{\sqrt{2x+a}}\\[0.3cm]
q=\frac{\partial z}{\partial y}=\frac{2}{2\sqrt{2y+b}}\equiv\frac{1}{\sqrt{2y+b}}\\[0.3cm]"

Then,


"\frac{1}{p}+\frac{1}{q}=\frac{1}{\frac{1}{\sqrt{2x+a}}}+\frac{1}{\frac{1}{\sqrt{2y+b}}}\longrightarrow\\[0.3cm]
\frac{1}{p}+\frac{1}{q}=\sqrt{2x+a}+\sqrt{2y+b}\equiv z\longrightarrow\\[0.3cm]
\boxed{z=\sqrt{2x+a}+\sqrt{2y+b}\longrightarrow z=\frac{1}{p}+\frac{1}{q}}"


Question (ii)



"z^2+\mu=2\left(1+\lambda^{-1}\right)\left(x+\lambda y\right)\longrightarrow\\[0.3cm]
2z\cdot\frac{\partial z}{\partial x}=2\left(1+\lambda^{-1}\right)\rightarrow p=\frac{1+\lambda^{-1}}{z}\\[0.3cm]
2z\cdot\frac{\partial z}{\partial y}=2\lambda\left(1+\lambda^{-1}\right)\rightarrow q=\frac{\lambda\left(1+\lambda^{-1}\right)}{z}\\[0.3cm]"



Then,



"\frac{1}{p}+\frac{1}{q}=\frac{1}{\frac{1+\lambda^{-1}}{z}}+\frac{1}{\frac{\lambda\left(1+\lambda^{-1}\right)}{z}}\longrightarrow\\[0.3cm]
\frac{1}{p}+\frac{1}{q}=\frac{z}{1+\lambda^{-1}}+\frac{z}{\lambda\left(1+\lambda^{-1}\right)}\longrightarrow\\[0.3cm]
\frac{1}{p}+\frac{1}{q}=\frac{z}{1+\frac{1}{\lambda}}+\frac{z}{\lambda+1}\longrightarrow\\[0.3cm]
\frac{1}{p}+\frac{1}{q}=\frac{z\lambda}{\lambda+1}+\frac{z}{\lambda+1}\longrightarrow\\[0.3cm]
\frac{1}{p}+\frac{1}{q}=\frac{z\left(\lambda+1\right)}{\lambda+1}\equiv z\\[0.3cm]
\boxed{z^2+\mu=2\left(1+\lambda^{-1}\right)\left(x+\lambda y\right)\longrightarrow z=\frac{1}{p}+\frac{1}{q}}"

Question 2


Let suppose that "b=b(a;\lambda,\mu)" . Then,



"F(x,y,z,a,b)=\sqrt{2x+a}+\sqrt{2y+b}-z=0\longrightarrow\\[0.3cm]
\frac{\partial F}{\partial a}=\frac{1}{2\sqrt{2x+a}}+\frac{1}{2\sqrt{2y+b}}\cdot\frac{db}{da}=0\longrightarrow\\[0.3cm]
\frac{1}{2\sqrt{2x+a}}=-\frac{1}{2\sqrt{2y+b}}\cdot\frac{db}{da}\longrightarrow\\[0.3cm]
2\sqrt{2y+b}=-2\sqrt{2x+a}\cdot\frac{db}{da}\longrightarrow\\[0.3cm]
\boxed{\sqrt{2y+b}=-\sqrt{2x+a}\cdot\frac{db}{da}}"


Let



"b(a)=-\frac{a}{\lambda}+C\longrightarrow\frac{db}{da}=-\frac{1}{\lambda}"

Then,



"\sqrt{2y+b}=-\sqrt{2x+a}\cdot\frac{db}{da}\longrightarrow\\[0.3cm]
\sqrt{2y+b}=\frac{1}{\lambda}\sqrt{2x+a}\longrightarrow\\[0.3cm]
\boxed{2y=\frac{2x+a}{\lambda^2}-b}\\[0.3cm]
\underbrace{\sqrt{2x+a}+\sqrt{2y+b}}_{z}=\sqrt{2x+a}-\sqrt{2x+a}\cdot\left(-\frac{1}{\lambda}\right)\\[0.3cm]
z=\left(1+\frac{1}{\lambda}\right)\sqrt{2x+a}\longrightarrow\\[0.3cm]
\boxed{z^2=\left(\frac{1+\lambda}{\lambda}\right)^2(2x+a)}"



Consider the found function along with



"z^2=2(1+\lambda^{-1})(x+y\lambda)-\mu\longrightarrow\\[0.3cm]
\boxed{z^2=2\left(\frac{1+\lambda}{\lambda}\right)(x+y\lambda)-\mu}"

Then,



"z^2=\left(\frac{1+\lambda}{\lambda}\right)^2(2x+a)\\[0.3cm]
z^2=2\left(\frac{1+\lambda}{\lambda}\right)(x+y\lambda)-\mu\\[0.3cm]
\left(\frac{1+\lambda}{\lambda}\right)^2(2x+a)=2\left(\frac{1+\lambda}{\lambda}\right)(x+y\lambda)-\mu\\[0.3cm]
\left(\frac{1+\lambda}{\lambda}\right)(2x+a)=2(x+y\lambda)-\frac{\lambda}{1+\lambda}\cdot\mu\\[0.3cm]
(1+\lambda)(2x+a)=2\lambda(x+y\lambda)-\frac{\lambda^2}{1+\lambda}\cdot\mu\\[0.3cm]
2x+a+2x\lambda+a\lambda=2x\lambda+2y\lambda^2-\frac{\lambda^2}{1+\lambda}\cdot\mu\\[0.3cm]
a(1+\lambda)=2y\lambda^2-2x-\frac{\lambda^2}{1+\lambda}\cdot\mu\\[0.3cm]
a(1+\lambda)=\left(\frac{2x+a}{\lambda^2}-b\right)\lambda^2-2x-\frac{\lambda^2}{1+\lambda}\cdot\mu\\[0.3cm]
a(1+\lambda)=2x+a-b\lambda^2-2x-\frac{\lambda^2}{1+\lambda}\cdot\mu\\[0.3cm]
b\lambda^2=-a\lambda-\frac{\lambda^2}{1+\lambda}\cdot\mu\\[0.3cm]
\boxed{b=-\frac{a}{\lambda}-\frac{\mu}{1+\lambda}}"



Conclusion,

The complete integral (ii) is the envelope of one parameter sub-system obtained by taking



"b=-\frac{a}{\lambda}-\frac{\mu}{1+\lambda}"